Nona is n years younger than Aida. 3n years later, Yati's age is twice of Nona's age. Aida is x years old now. Calculate the total age of Aida and Yati 3n years later.
Answer by ankor@dixie-net.com(22740)

Nona is n years younger than Aida.
N = A - n
3n years later, Yati's age is twice of Nona's age.
Y + 3n = 2(N + 3n)
Y + 3n = 2N + 6n
Y = 2N + 6n - 3n
Y = 2N + 3n
Aida is x years old now.
A = x
N = x - n
Y = 2(x-n) + 3n
Y = 2x - 2n + 3n
Y = 2x + n
Calculate the total age of Aida and Yati 3n years later.
T = total 3n years later
T = (A+3n) + (Y+3n)
T = A + Y + 6n
Replace A with x,
T = x + Y + 6n
replace Y with (2x+n)
T = x + (2x+n) + 6n
T = 3x + 7n, total of A & Y 3n years later